# SE(3)-Equivariant and Noise-Invariant 3D Motion Tracking in Medical Images

### v1

@@ Abstract
Rigid motion tracking is paramount in many medical imaging applications where movements need to be detected, corrected, or accounted for. Modern strategies rely on convolutional neural networks (CNN) and pose this problem as rigid registration. Yet, CNNs do not exploit natural symmetries in this task, as they are equivariant to translations (their outputs shift with their inputs) but not to rotations. Here we propose EquiTrack, the first method that uses recent steerable SE(3)-equivariant CNNs (E-CNN) for motion tracking. While steerable E-CNNs can extract corresponding features across different poses, testing them on noisy medical images reveals that they do not have enough learning capacity to learn noise invariance. Thus, we introduce a hybrid architecture that pairs a denoiser with an E-CNN to decouple the processing of anatomically irrelevant intensity features from the extraction of equivariant spatial features. Rigid transforms are then estimated in closed-form. EquiTrack outperforms state-of-the-art learning and optimisation methods for motion tracking in adult brain MRI and fetal MRI time series. Our code is available at github.com/BBillot/equitrack.
@@ 2.0.3 Landmark registration
Another registration paradigm is based on detecting landmarks that are used to estimate transforms (although not necessarily rigid) with classical optimisation techniques[26, 19] or closed-form algorithms[27, 28]. Such landmarks or features can be handcrafted[29], or rely on geometric features[30] or image intensities[31]. However, these strategies are sensitive to the quality of the extracted landmarks and consequently of the data itself.
@@ 2.0.4 Learning-based registration
Figure 1: Overview of EquiTrack. The fixed and moving volumes are first processed with a denoising CNN that removes anatomically irrelevant intensity features (noise, histogram shifts, etc.), so that its outputs only differ by the unknown rigid transform. Crucially, we then use a steerable SE(3)-equivariant E-CNN to extract matching anatomical features across images. A rigid transform $\hat{T}$ is estimated by computing summary statistics (centres of mass), providing us with two corresponding point clouds that are registered with a differentiable closed-form algorithm[28].
@@ 3.3.4 Steerable E-CNN
Steerable E-CNNs are formed by interleaving layers of the above-described filter banks with equivariant non-linearities, implemented as multiplicative equivariant scalar fields [13]. While steerable kernels have been derived in the continuous case, they can be easily discretised by sampling on $\mathbb{Z}^{3}$, followed by low-pass filtering with Gaussian smoothing to avoid aliasing [13]. During training, all $\kappa_{lj,J,m}$ remain fixed and we only learn the coefficients $\nu_{lj,J,m}$. Here, we implement $\Phi$ as a steerable E-CNN with scalar inputs (the images) and outputs (the $K$-channel feature maps).
@@ 3.4 Closed-form derivation of T
Multiple closed-form solutions have been proposed for (10)[55, 27, 28]. In order to put more emphasis on features with high activations, we introduce channel weights $\{w_{k}\}$ such that (10) now becomes:
@@
This formulation also has a closed-form solution [28]. Given a set of channel weights (described later in this section), the optimal translation $\hat{t}$ is estimated by subtracting the weighted centroids of the moving and fixed point clouds: $\hat{t}=\bar{x}^{m}-\bar{x}^{f}$ with $\bar{x}^{f}=\sum_{k}w_{k}x_{k}^{f}$ and $\bar{x}^{m}=\sum_{k}w_{k}x_{k}^{m}$. Next, the rotation $\hat{R}$ is derived as follows. We recenter all points around their centroid $\bar{x}^{f}_{k}=x^{f}_{k}-\bar{x}^{f}$ and stack them into matrices $\bar{X}^{f}$ and $\bar{X}^{m}$. If $W=\text{diag}(w_{1},...,w_{k})$, then $\hat{R}=VU^{T}$, where $U\Lambda V^{T}$ is the singular value decomposition of the weighted cross-correlation matrix $\Sigma=(\bar{X}^{m})^{T}W\bar{X}^{f}$.
Here, we define the channel weights as the product of image-specific weights, $w_{k}=\!w_{k}^{f}\!\times\!w_{k}^{m}$, where $\{w^{f}_{k}\}$ and $\{w^{m}_{k}\}$ are the positive sum of activations $\tilde{w}^{f}_{k}=\sum_{\Omega}|\phi_{k}^{f}|$ normalised across channels $w^{f}_{k}=\tilde{w}^{f}_{k}/\sum_{k}\tilde{w}^{f}_{k}$ (similarly for $\{w^{m}_{k}\}$).
@@ 3.5 Learning and training losses
EquiTrack is fully differentiable, yet we find that end-to-end training suffers from vanishing gradients through the E-CNN $\Phi$, making the training of $\Psi$ slow and unstable. We therefore train $\Psi$ and $\Phi$ separately and combine them at inference.
@@ 3.6.2 Architecture and training parameters
We implement the denoiser $\Psi$ as a 3D UNet [57] with 4 levels, each with 2 convolutional layers (32 $3\times 3\times 3$ kernels), ReLU non-linearities, and batch normalisation [58]. Inspired by [59, 60], we remove the top skip-connection, which we found to affect reconstructions by reintroducing early noisy features at the last stage of the UNet.
@@ 4.1.2 Fetal-I
We use a dataset of 55 whole-uterus 3D MRI time series with fetuses of gestational age from 25 to 35 weeks. Mothers are in a supine or lateral position. Scans are acquired on a 3T Skyra Siemens scanner using multi-slice single-shot gradient echo EPI sequences at 3mm isotropic resolution (TR = [5-8]ms, TE = [32-38]ms, $\alpha$ = ${}^{\circ}$). The mean image grid size is $120\!\times\!120\!\times\!80$, which we crop and pad to $96^{3}$. Intensities are rescaled to [0,1]. Automatic brain masking is applied [65] and masks are dilated by 4 voxels to account for segmentation uncertainty. We split the time series into 30 for training (N=764 scans in total), 5 for validation (N=152), and 20 for testing (N=633).
@@ 4.1.3 Fetal-II
We include a dataset of 23 clinical fetal MRI time series from 14 subjects. Importantly, all subjects are Chiari-II fetuses, characterised by cerebral and spinal malformations [66]. Further, these acquisitions use a gradient echo EPI sequence that is different from Fetal-I (TR=6.5s, TE=37ms, 2mm isotropic resolution). Volumes are preprocessed as before. These scans (N=614) are only used for testing, in order to evaluate the generalisation of our method.
Figure 2: Box plots for rotation errors, translation errors, and Dice scores on simulated pairs from the Adult and Fetal-I datasets. EquiTrack is significantly better than all other methods at the 5% level (Bonferroni-corrected two-sided non-parametric Wilcoxon signed-rank test), except for ANTs in terms of rotation error and Dice in adults (no statistically significant difference).
Figure 3: Sample registrations for all methods on pairs simulated from Fetal-I with small and large movements. The brain mask of the fixed image is shown in red for reference. Smoothness in registered images is due to interpolation. For small movements (top), all methods yield accurate or adequate results. In the case of large motion (bottom), only EquiTrack and ANTs (although less accurate) produce correct alignments.
@@ 4.2 Baseline methods
We test seven variants of four main competing methods. Every learning-based strategy is trained twice, and we select the model with the best validation scores.
@@ 4.2.1 ANTs [24]
This method rigidly registers scans by regressing the six transform parameters (one translation and rotation for each axis) using three convolution layers followed by three fully connected layers (FC). Volumes are concatenated as two channels when provided to the network, which is trained with a supervised geodesic loss on the transforms.
@@ 4.2.3 DLIR [33]
This pipeline is very similar to 3DPose-Net except that input volumes are processed separately by the CNN (like in EquiTrack) and the extracted features are concatenated when processed by the fully connected layers. We train two variants of this method: using the original L2 image loss (DLIR) or the same geodesic loss as 3DPose-Net (DLIR-geo). For fair comparison, we match the number of layers to that of 3DPose-Net, which slightly improves performance with respect to the original implementation.
@@ 4.2.4 KeyMorph [35]
KeyMorph is a state-of-the-art learning method for rigid registration. It learns to extract keypoints in an unsupervised fashion (i.e., using the L2 image loss) and aligns them with a closed-form algorithm. KeyMorph differs from EquiTrack in three aspects: a single CNN replaces the denoising CNN and E-CNN; this CNN is pre-trained with self-supervised learning; and transforms are estimated with a closed-form solution derived by taking the derivative of Eq. (10). We implement the original KeyMorph method, as well as two variants: one using our closed-form algorithm (KeyMorph-SVD) and one trained with a geodesic loss (KeyMorph-geo).
@@ 5.1 Testing on simulated pairs
The first experiment evaluates all methods on the adult and fetal MRI data. Here we use simulated pairs that provides us with exact ground truth transforms. These are created from anchor volumes that we augment twice with more realistic intensity and spatial transforms than for training: rotations in [${}^{\circ}$, ${}^{\circ}$], translations in [-6,6] vox., ${\sigma_{B}}_{\max}=0.2$, $\sigma_{\gamma}=0.2$, ${\sigma_{\xi}}_{\max}\!=\!0.03$ (these values are studied next in a sensitivity analysis). In total, we generate 200 pairs from Adult and 300 pairs from Fetal-I. We measure accuracy by averaging translation and rotation errors along the three axes, and by computing Dice between the fixed and moved brain masks.
@@
The results (Fig.2) reveal that EquiTrack accurately tracks 3D rigid motion in adult and fetal brain MRIs, with mean angle and translation errors below ${}^{\circ}$ and 0.5 voxels in all cases. Importantly, EquiTrack statistically outperforms all learning methods. While differences can be moderate on Adult (e.g., 2 Dice points with KeyMorph-SVD), the gap is considerably larger on Fetal-I (10 Dice points). This emphasises the difficulty of analysing noisy fetal data (Fig.3), necessitating models with enhanced robustness. This is the case of our architecture, where subdividing the registration task into removal of intensity changes and extraction of unbiased spatial features enables us to perform better than single CNNs, like KeyMorph. Meanwhile, we observe that processing images in parallel (i.e., DLIR vs. 3DPose-Net) and using an unsupervised training loss (as opposed to geodesic variants) are generally beneficial. Both concepts are implemented in EquiTrack. Finally, the fact that our closed-form algorithm accounts for potential landmark misalignments enables us to slightly improve the results of the original KeyMorph method: KeyMorph-SVD is better than KeyMorph by ${}^{\circ}$ and ${}^{\circ}$ on Adult and Fetal-I, respectively.
Table 1: Number of parameters and mean inference time on Fetal-I.
ANTs 3DPose-Net DLIR KeyMorph EquiTrack
[…]
Inference time (s) 7.3 0.10 0.11 0.37 0.51
Figure 4: Sensitivity study (mean and standard deviation) on Fetal-I for rotation, translation, bias field and noise. We start at translation=2, rotation=15, bias=noise=0; and increase the effect of each augmentation separately. Red indicates the maximum values used to simulate testing pairs (Fig.2).
@@
While ANTs and EquiTrack obtain similar results on the Adult dataset (no statistically significant difference in rotation error and Dice), our method consistently outperforms ANTs on Fetal-I, with mean improvements of 3.9 in Dice and ${}^{\circ}$ in rotation. Moreover, EquiTrack is much more robust than ANTs on Fetal-I: the edge of rotation and translation error boxes extend to ${}^{\circ}$ and 1.6 vox. for ANTs, as opposed to ${}^{\circ}$ and 0.3 vox for EquiTrack. Also, our method is an order of magnitude faster than ANTs (Table 1), which is critical for the deployment of tracking methods.

### v2

# SE(3)-Equivariant and Noise-Invariant 3D Rigid Motion Tracking in Brain MRI

@@ Abstract
Rigid motion tracking is paramount in many medical imaging applications where movements need to be detected, corrected, or accounted for. Modern strategies rely on convolutional neural networks (CNN) and pose this problem as rigid registration. Yet, CNNs do not exploit natural symmetries in this task, as they are equivariant to translations (their outputs shift with their inputs) but not to rotations. Here we propose EquiTrack, the first method that uses recent steerable SE(3)-equivariant CNNs (E-CNN) for motion tracking. While steerable E-CNNs can extract corresponding features across different poses, testing them on noisy medical images reveals that they do not have enough learning capacity to learn noise invariance. Thus, we introduce a hybrid architecture that pairs a denoiser with an E-CNN to decouple the processing of anatomically irrelevant intensity features from the extraction of equivariant spatial features. Rigid transforms are then estimated in closed-form. EquiTrack outperforms state-of-the-art learning and optimisation methods for motion tracking in adult brain MRI and fetal MRI time series. Our code is available at https://github.com/BBillot/EquiTrack.
@@ 2.0.3 Landmark registration
Another paradigm is based on detecting landmarks to estimate transforms (not necessarily rigid) with classical optimisation [26, 19] or in closed-form[27, 28]. Such landmarks can be handcrafted[29], or rely on intensities[30] or geometric features[31], such as second order derivatives, which have been applied to fetal motion correction [32]. However, these strategies are sensitive to the quality of the extracted landmarks and thus of the data itself.
@@ 2.0.4 Learning-based registration
Figure 1: Overview of EquiTrack. The fixed and moving volumes are first processed with a denoising CNN that removes anatomically irrelevant intensity features (noise, histogram shifts, etc.), so that its outputs only differ by the unknown rigid transform. Crucially, we then use a steerable SE(3)-equivariant E-CNN to extract $K$ matching anatomical features across images. A rigid transform $\hat{T}$ is estimated by computing summary statistics (centres of mass), providing us with two corresponding point clouds that are registered with a differentiable closed-form algorithm[28].
@@ 3.3.4 Steerable E-CNN
Steerable E-CNNs are formed by interleaving layers of the above-described filter banks with equivariant non-linearities, implemented as multiplicative equivariant scalar fields [13]. While steerable kernels have been derived in the continuous case, they can be easily discretised by sampling on $\mathbb{Z}^{3}$, followed by low-pass filtering with Gaussian smoothing to avoid aliasing [13]. During training, all filters $\kappa_{lj,J,m}$ remain fixed and we only learn the coefficients $\nu_{lj,J,m}$. Here, we implement $\Phi$ as a steerable E-CNN with scalar inputs (the images) and outputs (the $K$ feature maps).
@@ 3.4 Closed-form derivation of T
Multiple closed-form solutions have been proposed for (10)[56, 27, 28]. While (10) considers all $K$ extracted points equally, these points might not be predicted by the E-CNN with the same confidence. Hence, we introduce channel weights $\{w_{k}\}_{k=1}^{K}$ to account for the reliability of the corresponding points, and the previous equation now becomes:
@@
Here, we use the magnitude of the activations of $\Phi$ as a proxy for confidence: high activations indicate strong match between the extracted features and the representations learned by the network, while low activations denote weaker correlations that should be regarded as less reliable. Specifically, we define two sets of image-specific weights $\{w^{f}_{k}\}$ and $\{w^{m}_{k}\}$ as the sum of absolute activations $\tilde{w}^{f}_{k}=\sum_{\Omega}|\phi_{k}^{f}|$ normalised across channels $w^{f}_{k}=\tilde{w}^{f}_{k}/\sum_{k}\tilde{w}^{f}_{k}$ (similarly for $\{w^{m}_{k}\}$). We then obtain the final channel weights $\{w_{k}\}$ by combining the image-specific weights $w_{k}=\!w_{k}^{f}\!\times\!w_{k}^{m}$.
More generally, Eq. (11) also has a closed-form solution [28]. First, the optimal translation $\hat{t}$ is estimated by subtracting the weighted centroids of the moving and fixed point clouds: $\hat{t}=\bar{x}^{m}-\bar{x}^{f}$ with $\bar{x}^{f}=\sum_{k}w_{k}x_{k}^{f}$ and $\bar{x}^{m}=\sum_{k}w_{k}x_{k}^{m}$. The rotation $\hat{R}$ is then derived as follows. We recenter all points around their centroid $\bar{x}^{f}_{k}=x^{f}_{k}-\bar{x}^{f}$ and stack them into matrices $\bar{X}^{f}$ and $\bar{X}^{m}$. If $W=\text{diag}(w_{1},...,w_{k})$, then $\hat{R}=VU^{T}$, where $U\Lambda V^{T}$ is the singular value decomposition (SVD) of the weighted cross-correlation matrix $\Sigma=(\bar{X}^{m})^{T}W\bar{X}^{f}$.
@@ 3.5 Learning and training losses
EquiTrack is fully differentiable, yet end-to-end training suffers from vanishing gradients through $\Phi$ (due to the gated nonlinearities [13]), making the training of $\Psi$ unstable. Thus, we train $\Psi$ and $\Phi$ separately and combine them at inference.
@@ 3.6.2 Architecture and training parameters
We implement the denoiser $\Psi$ as a 3D UNet [58] with 4 levels, each with 2 convolution layers of 32 $3\!\times\!3\!\times\!$ 3 × 3 kernels, ReLU activations, and batch normalisation [59]. Inspired by [60, 61], we remove the top skip-connection, which affected reconstructions by reintroducing early noisy features at the last stage of the UNet.
@@ 4.1.2 Fetal-I
We use a dataset of 55 whole-uterus 3D MRI time series with fetuses of gestational age from 25 to 35 weeks. Mothers are in a supine or lateral position. Scans are acquired on a 3T Skyra Siemens scanner using multi-slice single-shot gradient echo EPI sequences at 3mm isotropic resolution (TR = [5-8]ms, TE = [32-38]ms, $\alpha$ = 90). The mean image grid size is $120\!\times\!120\!\times\!80$, which we crop and pad to $96^{3}$. Intensities are rescaled to [0,1]. Automatic brain masking is applied and masks are dilated by 4 voxels to account for segmentation uncertainty. We split the time series into 30 for training (N=764 scans in total), 5 for validation (N=152), and 20 for testing (N=633).
@@ 4.1.3 Fetal-II
We include a dataset of 23 clinical fetal MRI time series from 14 subjects. Importantly, all subjects are Chiari-II fetuses, characterised by cerebral and spinal malformations [66]. Further, these acquisitions use a gradient echo EPI sequence that is different from Fetal-I (TR=6.5s, TE=37ms, 2mm isotropic resolution). Volumes are preprocessed as before (see Appendix 8 for more details on the skull stripping step). These scans (N=614) are only used for testing, in order to evaluate the generalisation of our method.
Figure 2: Box plots for rotation errors, translation errors, and Dice scores on simulated pairs from the Adult and Fetal-I datasets. EquiTrack is significantly better than all other methods at the 5% level (Bonferroni-corrected two-sided non-parametric Wilcoxon signed-rank test), except for RPM and ANTs in terms of rotation error and Dice in adults (no statistically significant difference).
Figure 3: Example of the intermediate representations of EquiTrack. A reference scan (left) is either augmented with a large intensity transform (top row), a large spatial deformation (middle), or both (bottom). The denoiser $\Psi$ then removes noisy intensity features when appropriate (note how the middle example is left intact as no noise was added to it). After having removed intensity discrepancies across scans, the E-CNN $\Phi$ can now process volumes that only differ by their pose, which ensures the extraction of matching features across poses due to the SE(3)-equivariance of $\Phi$. These features are then used to accurately estimate rigid transforms back to the original volume whose brain mask is outlined in red for reference.
@@ 4.2 Baseline methods
We test seven variants of five competing baselines. We train each learning method twice, and select the model with the best validation scores.
4.2.1 Robust Point Matching (RPM) [67]
This rigid registration technique extracts surface meshes (e.g., with the marching cubes algorithm [68]) from segmentations of the objects to align. These meshes are then converted into continuous distributions using Gaussian Mixture Models, which are iteratively aligned at different scales with numerical optimisation (using a quasi-Newton algorithm). We select the parameters that maximise the validation scores: stride of 2 for marching cubes, and 3 levels of optimisation (scale parameter = 0.5, 0.2, 0.1).
@@ 4.2.2 ANTs [24]
This method rigidly registers scans by regressing the six transform parameters (one translation and rotation for each axis) using three convolution layers followed by a regression head of three fully connected layers. Volumes are concatenated before being fed to the network, which is trained with the same L2 image loss as EquiTrack (slightly better results than the original geodesic loss).
@@ 4.2.4 DLIR [34]
This strategy is similar to 3DPose-Net except that input scans are processed in parallel by the CNN (like in EquiTrack) and the extracted features are concatenated before the regression head. For a fair comparison, we match the number of layers to that of 3DPose-Net, which slightly improves the performance of the original implementation.
@@ 4.2.5 KeyMorph [36]
KeyMorph is a state-of-the-art learning method for rigid registration. It learns to extract keypoints in an unsupervised fashion using the L2 image loss and aligns them with a closed-form algorithm. KeyMorph differs from EquiTrack in three aspects: a single CNN replaces the denoising CNN and E-CNN; this CNN is pre-trained with self-supervised learning; and uses a different closed-form solution to estimate transforms, derived by taking the derivative of (10). We implement KeyMorph and a variant that uses our closed-form algorithm (KeyMorph-SVD). We also test another variant to study the effect of preprocessing KeyMorph’s inputs with the trained denoiser of EquiTrack (KeyMorph-denoiser).
Figure 4: Sample registrations for representative methods on 3D pairs simulated from Fetal-I with small and large movements. The brain mask of the fixed image is shown in red for reference. Smoothness in registered images is due to interpolation. For small movements (top), all methods yield accurate or adequate results. In the case of large motion (bottom), only EquiTrack and ANTs (although less accurate) produce correct alignments.
Table 1: Number of parameters for all methods (in order of magnitude), and training/inference times on Fetal-I.
The denoiser and E-CNN components of EquiTrack are trained separately.
RPM ANTs 3DPose-Net DLIR KeyMorph EquiTrack
Denoiser alone E-CNN alone Combined
Number of total parameters - - 2$\times$108 1$\times$108 6$\times$106 5$\times$106 2$\times$106 7$\times$106
Number of learnable parameters - - 2$\times$108 1$\times$108 6$\times$106 5$\times$106 8$\times$104 5$\times$106
Time per training iteration (s) - - 0.30 0.22 0.59 0.32 0.52 -
Inference time (s) 4.2 7.3 0.15 0.12 0.37 0.17 0.34 0.51
Figure 5: Example of features produced by KeyMorph-SVD and EquiTrack on a simulated pair from Fetal-I. The regular CNN of KM can lead to inconsistent representations across poses (red arrows). In contrast, The E-CNN used in EquiTrack guarantees the extraction of SE(3)-equivariant features, which enable accurate rigid registration.
Figure 6: Sensitivity study (mean $\pm$ standard deviation) on Fetal-I for rotation, translation, bias field and noise. We start at translation =\!=\!= 2 vox., rotation =\!=\!= 15, bias =\!=\!= noise =\!=\!= 0; and increase the effect of each augmentation separately. Red denotes the maximum values used to simulate testing pairs in Fig.2. E-CNN alone is a version of EquiTrack without a denoiser (the two methods have overlapping curves for the translation and rotation studies). In contrast to the baselines, EquiTrack is robust to both large motion (E-CNN) and intensity variations (denoising CNN) at the same time.
@@ 5.1 Testing on simulated pairs
The first experiment evaluates all methods on the adult and fetal MRI data. Here we use simulated pairs that provides us with exact ground truth transforms. These are created from anchor volumes that we augment twice with more realistic intensity and spatial transforms than for training: rotations in [-45, 45], translations in [-6,6] vox., ${\sigma_{B}}_{\max}=0.2$, $\sigma_{\gamma}=0.2$, ${\sigma_{\xi}}_{\max}\!=\!0.03$ (these values are studied next in a sensitivity analysis). In total, we generate 200 pairs from Adult and 300 pairs from Fetal-I. We measure accuracy by averaging translation and rotation errors along the three axes, and by computing Dice between the fixed and moved brain masks.
@@
The results (Fig.2) reveal that EquiTrack accurately tracks 3D rigid motion in adult and fetal brain MRIs, with mean angle and translation errors below 4 and 0.5 voxels in all cases. Importantly, EquiTrack statistically outperforms all learning methods. While differences can be moderate on Adult (e.g., 2 Dice points with KeyMorph-SVD), the gap is considerably larger on Fetal-I with 10 Dice points. This emphasises the difficulty of analysing noisy fetal data, necessitating models with enhanced robustness. This is the case of our architecture, where using an E-CNN guarantees the extraction of matching features across poses (Fig. 3). Indeed, in contrast with regular CNNs, the E-CNN used in EquiTrack enables consistent latent representations (Fig. 5), which improves the quality of motion estimates compared to the other learning methods (Fig. 4). Meanwhile, we observe that processing images in parallel (i.e., DLIR vs. 3DPose-Net) is generally beneficial and is the solution implemented in EquiTrack. Interestingly, KeyMorph and KeyMorph-denoiser yield almost identical scores, showing that regular CNNs (as opposed to E-CNNs) do not need a denoiser to model noise invariance when trained with appropriate intensity augmentations. Finally, the fact that our closed-form algorithm accounts for potential landmark misalignments enables us to slightly improve the results of the original KeyMorph method: KeyMorph-SVD is better than KeyMorph by 1.7 and 2.3 on Adult and Fetal-I, respectively.
@@
While the optimisation methods (RPM and ANTs) obtain results similar to EquiTrack on the Adult dataset (no statistical difference in rotation error and Dice), our method consistently outperforms them on Fetal-I by at least 2.5 in rotation and 3.9 Dice points. We note that ANTs is slightly better than RPM for the fetal dataset, which highlights the benefits of exploiting intracranial intensity features in the images (ANTs), as opposed to purely relying on the segmentation mask (RPM). Importantly, EquiTrack is much more robust than ANTs on Fetal-I: the edge of rotation and translation error boxes extend to 30.6 and 1.6 vox. for ANTs, as opposed to 4.6 and 0.3 vox for EquiTrack. Also, our method is an order of magnitude faster than the optimisation methods (Table 1), which is critical for online deployment of tracking methods.